of the studnts who play in the band, 2/3 are eighth graders, 1/6 are seventh graders and 5 sixth graders. How many students are in the band?
Answer by edjones(8007)

Let x=total students in the band.
2x/3 + x/6 + 5 = x
4x+x+30=6x Multiply each side by 6 to eliminate fractions.
6x=5x+30
x=30 students
.
Check:
60/3 + 30/6 + 5 = 30 true
.
Ed